Prev Next

SW&IT Meeting Minutes 2022-01-24

Attendees: Brandon, David, Glenn, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20220124
LR20220110 - Bec, Matthieu 72.0
LR20220110 - Mechtley, Brandon 80.0
LR20220110 - Eychaner, Glenn 72.0
LR20220110 - Edmeades, David 72.0
LR20220110 - Grenz, Paul 36.0
LR20220110 - Hooper, Stephen 0.0
LR20220110 - Zhang, Yang 72.0
LR20220110 - Zhang, Xianyu 72.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

  • network outage, actions and organization

Team reports

MB

  • Last week
    • SOUL/UAO SW installation discussion
    • LN planning
    • iLocater monthly
    • LBTI archive scheme
    • DMS repository / lmircam
    • DMS lit2/webpack
  • This week
    • ATT PM
    • ARGOS / UAO
    • DMS archive index

YZ

  • Previous:
    • IT8315, updated GCS and released and installed a patch (no word yet on whether the problem is fixed!).
    • lib-telemetry debug (for the leap seconds complaints). Found a bug that caused the trouble!
    • Some MODS image cleaning troubles (on-going discussion with Olga).
    • TMS meeting and software work.
    • Some issuetraks.
  • Next:
    • Update lib-telemetry and maybe install the new version.
    • Release a new version of TMS software with missing channel handling (plus other TMS work).
    • Work on MODS image cleaning with Olga.
    • Maybe IT7400?

DE

  • JW laptop setup, waiting on input from him re: software packages
  • Vault work
    • New Vault set up, early tests good.
    • Cudgeling Ketiv into understanding that we want to change the IIS config to get Vault thin client on the webroot
    • Modified directory permissions on the NAS to allow unpriviliged users to write to backup share, which allowed me to modify the backup script and
    • I should be able to remove and not replicate the large staging volume that was taking up a significant amount of local VDisk storage
    • Anticipate being able to announce to Engineering that new Vault is open for testing, after which
    • I will announce the cutover generally. Ketiv says that it should handle checked-out files but it would be safer to ask everyone to check everything in.
  • Veeam licensing shakeup; getting info on costs. I strongly suspect that we will be better served by preemptively purchasing enough current-style socket licenses to cover our expected future needs.
    • The new style licenses are more flexible but I can only imagine that they're doing the change to get more money.
    • I'll report back when I get a response from CDW.
  • Archive newdata
    • Have an iSCSI device set up on mountain Synology, mounted on archive-in, and ready to swap. I'll need some downtime to reboot it and make sure that all the resources and services are persistent and to do the swap.
    • Current size is 10TiB.
    • Not particularly difficult to make larger, but is enough steps that we should not rely on doing that in an emergency over the phone.
    • Once we expand, we can't shrink the volume, so we'd have to clear it and recreate it if shrinking were needed.
  • Looks like we'll get the RMA SSD back from Samsung faster than the bought one. I'm not sure that that exists at all.
  • VPN account make, I am assured that they'll test it before their runs. We shall see.

GE

  • Last week
    • Finished the changes to the RTD and LUCI software needed for subwindowing and deployed to obs5.
    • Merged the AO fix (IT 8530) and deployed to luci.luci.
  • This week
    • Performance testing of the new RTD; see the notes in the pull request for details.
    • Merge the new RTD after D.Thompson finishes testing, and deploy to the observer computers.
    • Begin work on the observer GUI; IT 6138 (problem with pauses in GUI) and IT 8529 (GUI possibly disconnecting from luci.luci).
    • Examine the other LUCI GUIs to check for GUI modifications from threads other than the Event Dispatch Thread.

BM

  • Last week:
    • AO-SW meeting, LN planning meeting, AOG meeting
    • Set up ao sw on DX soul, adsec machines
    • LBTI issues:
      • Offloading (IT 8556): tried with LBTI team to reproduce last month, Amali has ideas for reproducing–need to know when this will happen.
      • ICE communication errors: discussed with Xianyu and Juan Carlos.
      • Button: Still need to test with both a WFS and Adsec simulator. Need to set up WFS VM.
  • This week:
    • ARGOS SW meeting Wednesday.
    • OCAM tests: pending PR from Fabio.
    • LBTI issues:
      • Figure out what, if any tests will take place to reproduce the problem.
      • ICE communication error fix. Thanks Xianyu.
      • Manual SafeSkip button testing.
    • Make PRs for changes from DEC night–wrong filotto logging, SafeSkip logging
    • Catch up inspecting logs for adsec failures from last week.

PSG

  • Last week
    • Testing revealed the need to add "reset" to OSS GUI for ARGOS swing arms.
    • Added new functionality to swing arm - "reset".
    • Completed "Conflict of Interest" training.
  • This week
    • Schedule and test new functionality.
    • Possibly adding ARGOS defaults to OSS.conf file (used during reset - now hard coded in web page).
    • Continue looking at GCS code.

SH

  • New Synology network, cluster formation
    • Wiki on the new Synology
  • New core switches
  • VM2 Tucson
  • Inventory with Louise
  • URIC UPS battery (finally)

next:

  • Switch more data over to the new Synology
    • archive and repository have been copied
    • CV's data has been copied.
    • Home directories with Google LDAP shoule be interesting.
  • VMware work in Tucson
Topic revision: r11 - 24 Jan 2022, GlennEychaner
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback